How is this questions solved?
Consider the polynomial function p(x)=2x^3-9x^2+4x+15
Show using the intermediate therem that p(x) has a zero between -2 and 0.

Answer by nerdybill(7384)

See the web site below for additional explanation:
http://www.mathsisfun.com/algebra/intermediate-value-theorem.html
.
p(x)=2x^3-9x^2+4x+15
.
p(-2)=2(-2)^3-9(-2)^2+4(-2)+15
p(-2)=2(4)-9(4)+(-8)+15
p(-2)=8-36-8+15
p(-2)=-36+15
p(-2)=-21 (negative)
.
p(0)=2(0)^3-9(0)^2+4(0)+15
p(0)=0-0+0+15
p(0)=15
.
Since p(-2) is NEGATIVE the function MUST cross y=0 to get to p(0). Because p(0) is POSITIVE.